OpenStudy (anonymous):

√0.49= -7 0.07 0.7 7

OpenStudy (anonymous):

0.7

OpenStudy (anonymous):

0.7

OpenStudy (anonymous):

0.7

OpenStudy (anonymous):

0.7

OpenStudy (dumbcow):

\[\sqrt{0.49} = \sqrt{49/100} = \frac{\sqrt{49}}{\sqrt{100}} = \frac{7}{10} = 0.7\]

OpenStudy (anonymous):

\[\sqrt{\frac{49}{100}}=\frac{7}{10}\]